# Model-oriented Graph Distances via Partially Ordered Sets

### v1

@@ 3.3.3 Relation to edge-by-edge distances
Recall that SHDs measure the distance between two graphs by counting how many edges are mismatched. They are therefore by construction edge-by-edge style distance, in the sense that they can be computed considering each edge at a time. Different types of mismatches are often assigned different weights (Perrier, Imoto and Miyano, 2008). For PDAGs, for example, there are two common variants of SHD (Tsamardinos, Brown and Aliferis, 2006) defined as follows. Consider encoding a PDAG $\mathcal{G}$ over $\mathrm{V}$ with adjacency matrix
@@
$\begin{aligned} \operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\text{ or }A^{s}_{j,i}\neq A^{t}_{j,i}\right],\\ \operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\right]+\mathbb{I}\left[A^{s}_{j,i}\neq A^{t}_{j,i}\right].\end{aligned}$
Here, $\mathbb{I}[\cdot]$ is the indicator function. In the following corollary to Theorem˜2, we study for general graph-types when the model-oriented distance also behaves edge-by-edge, and discuss the implications for certain PDAGs considered in this work.
Lemma 1.
Suppose we consider graphs $\mathfrak{G}$ with $k$ edge types (see Section 2.1.1) that define a poset $\mathcal{L}$, in which they are ordered based on subgraph containment; that is, $\mathcal{G}_{1}=(V,E^{1}_{1}\cup\dots\cup E_{1}^{k})\preceq\mathcal{G}_{2}=(V,E^{1}_{2}\cup\dots\cup E^{k}_{2})$ if $E^{i}_{1}\subseteq E^{i}_{2}$ for all $i$. Further, suppose $\mathcal{L}$ is graded with $\mathrm{rank}(\mathcal{G})=\sum_{i=1}^{k}|E_{i}|$ for any $\mathcal{G}\in\mathfrak{G}$, and that it is lower semimodular. Then, $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$, where $\triangle$ denotes the symmetric set difference. If the graph class contains only undirected edges, this is equal to $\operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})$. If the class contains only directed edges, this is equal to $\operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})$.
@@
Since the poset is lower semimodular, we have from Theorem 2 that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. By construction, $\mathcal{L}$ is a meet semi-lattice with $\mathcal{G}_{1}\wedge\mathcal{G}_{2}=(V,E_{1}^{1}\cap E_{2}^{1}\cup\dots\cup E_{1}^{k}\cap E_{2}^{k})$. Appealing to Corollary 1, we then have that: $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\mathrm{rank}(\mathcal{G}_{s})+\mathrm{rank}(\mathcal{G}_{t})-2\mathrm{rank}(\mathcal{G}_{s}\wedge\mathcal{G}_{t})=\sum_{i=1}^{k}|E^{1}_{i}\triangle E^{2}_{i}|$. ∎
Effectively, this shows that the model-oriented distance is an edge-by-edge distance whenever the partial ordering can be determined by examining one edge at a time (subgraph containment) and the Hasse diagram is sufficiently dense (lower semimodularity). We will see that this typically holds for graph classes that impose few constraints, such as UGs and DAGs, but not for CPDAGs or MPDAGs. Which version of the SHD it coincides with then depends on whether the edges are directed or undirected. In particular, as we will see in the next section, for UGs with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$, the model-oriented distance is equivalent to $\operatorname{\mathrm{SHD}}_{1}$. However, the model-oriented distance diverges from $\operatorname{\mathrm{SHD}}_{1}$ for all other graph types considered in this paper.
4 Model-oriented distance $d_{\mathcal{L}}$ in probabilistic and causal graphs
Next, we apply the poset framework described in Section˜3 to derive the model-oriented distance for UGs, DAGs, CPDAGs, and MPDAGs, presented in order of increasing complexity in their associated poset. Table˜1 highlights the different attributes for each associated poset, with the class of UGs being the most structured, and the class of MPDAGs being the least structured. We will formally show these properties in this section. For each graph class, we first study the corresponding model-oriented poset, then use it to derive the model-oriented distance, and finally examine simpler characterizations of the model-oriented distance, for example in terms of the SHD, or when this is not possible approximations to the model-oriented distance.
Table 1: Structural attributes of the underlying model-oriented poset for different graph classes
@@ 4 Model-oriented distance $d_{\mathcal{L}}$ in probabilistic and causal graphs
Polytree MPDAG causal empty graph
@@ 4.1 Probabilistic UGs
As a warm-up, consider $\mathfrak{G}$ the class of UGs over a finite vertex set $\mathrm{V}$ representing probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ given by Eq.˜1. The next result states that the model inclusion becomes graph inclusion, which we prove in Section˜A.2.
Proposition 4.
@@
From these properties, we obtain a closed form for the model-oriented distance.
@@ 4.2 Causal DAGs
Consider $\mathfrak{G}$ the class of DAGs over a finite vertex set representing causal models in the sense of $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ given by Eq.˜4. The partial order is given by subgraph containment, stated as follows.
Proposition 6.
@@
We prove Proposition˜6 in Section˜A.2. Using this result, the model-oriented poset can be characterized as follows; see Section˜E.1 for its proof. Fig.˜3(b) shows the Hasse diagram of this poset when the vertex set is of size 3.
Proposition 7.
@@
    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$.
@@
From Propositions˜7 and 1, we obtain the model-oriented distance as follows.
@@ 4.3 Probabilistic CPDAGs
Let $\mathfrak{G}$ be the class of CPDAGs over a finite vertex set $\mathrm{V}$, which represent probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$ given by Eq.˜2. We use $\mathcal{G}]$ to denote the set of Markov equivalent DAGs represented by a CPDAG $\mathcal{G}$. The model containment is characterized by a basic result due to Chickering (2002), restated as follows.
Lemma 2 (Chickering, 2002, Theorem 4).
@@
Here, an edge $u\rightarrow v$ is covered in DAG $\mathcal{D}$ if $\operatorname{Pa}_{\mathcal{D}}(v)=\operatorname{Pa}_{\mathcal{D}}(u)\cup\{u\}$. Using the result above, we can derive the following properties of CPDAG’s model-oriented poset. Fig.˜3(a) shows the Hasse diagram of this poset when $|\mathrm{V}|=3$.
Proposition 8.
@@
Additionally, $\mathcal{L}$ is neither a (join or meet) semi-lattice nor (upper or lower) semimodular in general. For proofs of these results, see Section˜E.2.
@@ 4.3.1 Upper and lower bounds on $d_{\mathcal{L}}$
The lack of semi-lattice and semimodularity structures makes computing our model-oriented distance between CPDAGs much harder than UGs and DAGs. First, Lemma˜1 is no longer applicable and indeed, as we demonstrate in Example˜1, the SHD behaves very differently from our distance. Further, by Theorem˜2, because the poset is not semimodular, up-down and down-up distances can only provide upper bounds on the model-oriented distance. Indeed, consider the example in Fig.˜5, where the shortest path, taking a ‘zigzag’ shape, is of length 4, while both the shortest down-up and up-down paths have length 8.
Figure 5: The model-oriented distance between probabilistic CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is 4, as shown by the shortest path in red. Meanwhile, the up-down distance is 8 (blue path) and the down-up distance is also 8 (orange path).
Due to these difficulties, in Section˜5, we propose using the $\text{A}^{\ast}$ algorithm to compute $d_{\mathcal{L}}$. As an informed search algorithm, $\text{A}^{\ast}$ relies on good, custom-made lower and upper bounds on the distance to speed up the search: the lower bound, often called the heuristic function in the literature, allows one to disregard some paths without discarding a potentially optimal path; the upper bound let one prune all the paths that are longer than the current best solution. In this section, we investigate lower and upper bounds on $d_{\mathcal{L}}$ between CPDAGs.
To formulate the lower bound we need some notation. For a PDAG $\mathcal{G}$, we use $\operatorname{sk}(\mathcal{G})$ to denote its skeleton, which is the set of edges after undirecting all the edges in $\mathcal{G}$. We also use $\operatorname{v-str}(\mathcal{G})$ to denote the set of triplets $(a,b,c)$ that form a v-structure at $b$ in $\mathcal{G}$ (see Section˜2.1.4); here we do not distinguish between $(a,b,c)$ and $(c,b,a)$. Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$. Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$. Define the set
@@
which consists of v-structures that appear in only one of the two graphs. The set above can always be partitioned as
@@
where two triplets $(a,b,c)$ and $(d,e,f)$ are placed in the same partition if they have any vertex in common (e.g., $a=d$ or $a=e$).
For a pair of vertices $(i,j)$, we define operations $\operatorname{OP}_{(i,j),\text{CE}}$ (cover edge) and $\operatorname{OP}_{(i,j),\text{ER}}$ (edge removal) that map a subset of $\mathrm{S}$ to a subset of $\mathrm{S}$, given by
$\begin{array}[]{ll}\operatorname{OP}_{(i,j),\text{CE}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,k,j):k\in\mathrm{V}\},\\ \operatorname{OP}_{(i,j),\text{ER}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,j,k),(j,i,k):k\in\mathrm{V}\}\end{array},\quad\mathrm{T}\subseteq\mathrm{S}.$
@@
$m_{\text{op}}(\mathrm{S}_{l}):=\min\left\{m\in\mathbb{N}:\exists~\mathrm{op}_{1},\dots,\mathrm{op}_{m}\text{ such that }\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S}_{i})=\emptyset\right\},\quad l=1,\dots,L,$
where each $\mathrm{op}_{k}$ is either $\operatorname{OP}_{(i,j),\text{CE}}$ or $\operatorname{OP}_{(i,j),\text{ER}}$ for some $(i,j)$ that can depend on $k$. Because the v-structures from two different partitions share no common vertex by construction and this property is maintained by both operations, we conclude that
$m_{\text{op}}(\mathrm{S}):=\sum_{l=1}^{L}m_{\text{op}}(\mathrm{S}_{l})$
lower bounds the number of steps taken to align the set of v-structures between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
Property 1 (Upper bound).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})\right\}.$
Property 2 (Lower bound).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2m_{\text{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
We prove Property 2 in Appendix F.1. Property 1 follows from recalling that the model-oriented distance is upper bounded by both the down-up and the up-down distance (both exist for CPDAGs since the poset has a least and greatest element), and appealing to Proposition 3.

### v2

@@ 3.3.3 Relation to edge-by-edge distances
Recall that the structural Hamming distance is given by the number of edges that are mismatched between the two graphs. Different types of mismatches can receive different weights (Perrier, Imoto and Miyano, 2008). For PDAGs, for example, there are two common variants of SHD (Tsamardinos, Brown and Aliferis, 2006) defined as follows. Consider encoding a PDAG $\mathcal{G}$ over $\mathrm{V}$ with adjacency matrix
@@
$\begin{aligned} \operatorname{\mathrm{SHD}}_{1}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i<j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\text{ or }A^{s}_{j,i}\neq A^{t}_{j,i}\right],\\ \operatorname{\mathrm{SHD}}_{2}(\mathcal{G}_{s},\mathcal{G}_{t})&:=\sum_{i,j}\mathbb{I}\left[A^{s}_{i,j}\neq A^{t}_{i,j}\right].\end{aligned}$
Here $\mathbb{I}[\cdot]$ is the indicator function. Note that $\operatorname{\mathrm{SHD}}_{2}\geq\operatorname{\mathrm{SHD}}_{1}$ holds by definition. The following result outlines when our distance reduces to an edge-by-edge type.
Lemma 1 (Edge-by-edge model oriented distances).
Let $\mathfrak{G}$ be a class of graphs with $k\geq 1$ edge types (see Section˜2.2.1). Consider a poset $\mathcal{L}$ that orders graphs by subgraph containment, i.e., $\mathcal{G}_{s}=(\mathrm{V},\mathrm{E}^{1}_{s},\dots,\mathrm{E}_{s}^{k})\preceq\mathcal{G}_{t}=(\mathrm{V},\mathrm{E}^{1}_{t},\dots,\mathrm{E}^{k}_{t})$ if $\mathrm{E}^{i}_{s}\subseteq\mathrm{E}^{i}_{t}$ for $i=1,\dots,k$. Suppose $\mathcal{L}$ is lower semimodular and graded with $\operatorname{\mathrm{rank}}(\mathcal{G})=\sum_{i=1}^{k}|\mathrm{E}^{i}|$ for each $\mathcal{G}=(\mathrm{V},\mathrm{E}^{1},\dots,\mathrm{E}^{2})$. Then, we have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\sum_{i=1}^{k}|\mathrm{E}^{i}_{s}\triangle\mathrm{E}^{i}_{t}|$. Further, when $k=1$, $d_{\mathcal{L}}$ reduces to $\operatorname{\mathrm{SHD}}_{1}$ if $\mathfrak{G}$ is a class of undirected graphs, and to $\operatorname{\mathrm{SHD}}_{2}$ if $\mathfrak{G}$ is a class of directed graphs.
@@
Since the poset is lower semimodular, we have from Theorem˜2 that $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t})$. By construction, $\mathcal{L}$ is a meet semi-lattice with $\mathcal{G}_{s}\wedge\mathcal{G}_{t}=(\mathrm{V},\mathrm{E}_{s}^{1}\cap\mathrm{E}_{t}^{1},\dots,\mathrm{E}_{s}^{k}\cap\mathrm{E}_{t}^{k})$. Appealing to Corollary˜1, we then have $d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})=\operatorname{\mathrm{rank}}(\mathcal{G}_{s})+\operatorname{\mathrm{rank}}(\mathcal{G}_{t})-2\operatorname{\mathrm{rank}}(\mathcal{G}_{s}\wedge\mathcal{G}_{t})=\sum_{i=1}^{k}\left(|\mathrm{E}^{i}_{s}|+|\mathrm{E}^{i}_{t}|-2|\mathrm{E}^{i}_{s}\cap\mathrm{E}^{i}_{t}|\right)=\sum_{i=1}^{k}|\mathrm{E}^{i}_{s}\triangle\mathrm{E}^{i}_{t}|$. ∎
Effectively, this shows that the model-oriented distance is an edge-by-edge distance whenever the partial ordering can be determined by examining one edge at a time (subgraph containment) and the Hasse diagram is sufficiently dense (lower semimodular). This typically holds for graph classes that impose few constraints, such as UGs and DAGs, but not for CPDAGs or MPDAGs. The version of the SHD it coincides with depends on whether the edges are directed or undirected. In particular, as we will see in the next section, for UGs with model map $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$, the model-oriented distance is equivalent to $\operatorname{\mathrm{SHD}}_{1}$. The model-oriented distance diverges from $\operatorname{\mathrm{SHD}}_{1}$ for all other graph types considered in this paper. We will compare our distance with SHD numerically in Section˜5.
4 Model-oriented distance for probabilistic and causal graphs
We are ready to apply the general poset framework to UGs, DAGs, CPDAGs, MPDAGs as well as simpler subclasses of CPDAGs and MPDAGs. Table˜2 summarizes the structural properties of the poset underlying each graph class, with UGs being the most structured and MPDAGs being the least structured. We will use this section to prove these properties and consider their implication on computing the model-oriented distance.
Table 2: Structural properties of the underlying model-oriented poset for different graph classes
@@ 4 Model-oriented distance for probabilistic and causal graphs
Polytree MPDAG causal empty graph lower
@@ 4.1 Probabilistic UGs
As a warm-up, consider $\mathfrak{G}$ the class of UGs over a finite vertex set $\mathrm{V}$ representing probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{ug}}$ given by Eq.˜1. The next result states that model inclusion becomes graph inclusion for UGs, which we prove in Section˜A.2.
Proposition 4 (UG comparability).
@@
From Theorems˜3 and 1, we obtain a closed form for the model-oriented distance.
@@ 4.2 Causal DAGs
Let $\mathfrak{G}$ be the class of DAGs over a finite vertex set representing causal models in the sense of $\mathcal{M}^{\mathrm{c}}_{\textsc{dag}}$ given by Eq.˜4. The partial order is given by subgraph containment, stated as follows.
Proposition 5 (DAG comparability).
@@
We prove Proposition˜5 in Section˜A.2. Using this result, the model-oriented poset is characterized by the next result, which is proved in Section˜E.1. Fig.˜3(b) shows the Hasse diagram of this poset when the vertex set is of size 3.
Theorem 4 (DAG poset).
@@
    $\mathcal{G}_{\hat{0}}$ is the empty graph. There is no $\mathcal{G}_{\hat{1}}$ when $|\mathrm{V}|>1$.
@@
From Theorems˜4 and 1, we obtain a closed form for the model-oriented distance as follows.
@@ 4.3 Probabilistic CPDAGs
Let $\mathfrak{G}$ be the class of CPDAGs over a finite vertex set $\mathrm{V}$, which represent probabilistic models in the sense of $\mathcal{M}^{\mathrm{o}}_{\textsc{cpdag}}$ given by Eq.˜2. For $\mathcal{G}\in\mathfrak{G}$, we use $\mathcal{G}]$ to denote the set of Markov equivalent DAGs it represents. The model containment is characterized by a basic result due to Chickering (2002), restated as follows.
Lemma 2 (CPDAG comparability).
@@
Here, a directed edge $u\rightarrow v$ is covered in DAG $\mathcal{D}$ if $\operatorname{Pa}_{\mathcal{D}}(v)=\operatorname{Pa}_{\mathcal{D}}(u)\cup\{u\}$. Using the result above, we can derive the properties of CPDAG’s model-oriented poset listed in Theorem˜5. Fig.˜3(a) shows the Hasse diagram of this poset when $|\mathrm{V}|=3$.
Remark 2.
In light of the second result in Lemma˜2, it may be tempting to conjecture that for CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$, we have $\mathcal{G}_{s}\preceq\mathcal{G}_{t}$ if and only if there exist DAGs $\mathcal{D}_{s}\in[\mathcal{G}_{s}],\mathcal{D}_{t}\in[\mathcal{G}_{t}]$ such that $\mathcal{D}_{s}$ is a subgraph of $\mathcal{D}_{t}$. But the ‘only if’ part of this is not true; consider the following two graphs over $\{1,2,3,4\}$:
$\mathcal{G}_{s}:\begin{array}[]{ccc}1&&2\\ &\searrow&\downarrow\\ 4&&3\end{array},\qquad\mathcal{G}_{t}:\begin{array}[]{ccc}1&\rightarrow&2\\ \uparrow&\nwarrow&\uparrow\\ 4&&3\end{array}.$
Taeb, Bühlmann and Chandrasekaran (2024) highlighted some properties of the CPDAG poset, which we restate as follows.
Theorem 5 (CPDAG poset).
@@
Additionally, $\mathcal{L}$ is neither a (join or meet) semi-lattice nor (upper or lower) semimodular in general. For completeness, we prove these results in Section˜E.2.
@@ 4.3.1 Upper and lower bounds on $d_{\mathcal{L}}$
The lack of semi-lattice and semimodularity structures makes computing our model-oriented distance between CPDAGs harder than for UGs and DAGs. First, Lemma˜1 is no longer applicable and indeed, as we demonstrate in Example˜1, the SHD behaves very differently from our distance. Further, by Theorem˜2, because the poset is not semimodular, up-down and down-up distances can only provide upper bounds on the model-oriented distance. Indeed, consider the example in Fig.˜5, where the shortest path, taking a zigzag shape, is of length 4, while both the shortest down-up and up-down paths have length 8.
Figure 5: The model-oriented distance between probabilistic CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ is 4, as shown by the shortest path drawn in red. Meanwhile, the up-down (blue path) and the down-up (orange path) distances are both 8.
Due to these difficulties, in Section˜5, we propose using the A* algorithm to compute $d_{\mathcal{L}}$. As an informed search algorithm, A* relies on good, custom-made lower and upper bounds on the distance to speed up the search: the lower bound, often called the heuristic function in the literature, allows one to disregard some paths without discarding a potentially optimal path; the upper bound enables early termination and further pruning of suboptimal paths. In this section, we investigate lower and upper bounds on $d_{\mathcal{L}}$ between CPDAGs.
To formulate the lower bound, we need some notation. For a PDAG $\mathcal{G}$, we use $\operatorname{sk}(\mathcal{G})$ to denote its skeleton, which is the set of edges upon undirecting all the edges in $\mathcal{G}$. We also use $\operatorname{v-str}(\mathcal{G})$ to denote the set of triplets $(a,b,c)$ that form a v-structure at $b$ in $\mathcal{G}$ (see Section˜2.2.4 for its definition); here we do not distinguish between $(a,b,c)$ and $(c,b,a)$. Let us fix two CPDAGs $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$ over a vertex set $\mathrm{V}$. Define the set
@@
which consists of v-structures that appear in only one of the two graphs. We partition $\mathrm{S}$ as
@@
where two triplets $(a,b,c)$ and $(d,e,f)$ are placed in the same partition if they share any common vertex (e.g., $a=d$ or $a=e$).
For a pair of vertices $(i,j)$, we define operations $\operatorname{OP}_{(i,j),\text{CE}}$ (cover edge) and $\operatorname{OP}_{(i,j),\text{ER}}$ (edge removal) that map a subset of $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$ to a subset of $\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})$, given by
$\begin{array}[]{ll}\operatorname{OP}_{(i,j),\text{CE}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,k,j):k\in\mathrm{V}\},\\ \operatorname{OP}_{(i,j),\text{ER}}(\mathrm{T})&:=\mathrm{T}\setminus\{(i,j,k),(j,i,k):k\in\mathrm{V}\}\end{array},\quad\mathrm{T}\subseteq\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}).$
@@
$m_{\text{op}}(\mathrm{S}_{l}):=\min\left\{m\in\mathbb{N}:\exists~\mathrm{op}_{1},\dots,\mathrm{op}_{m},\;\mathrm{op}_{1}\circ\dots\circ\mathrm{op}_{m}(\mathrm{S}_{l})=\emptyset\right\},\quad l=1,\dots,L,$ (6)
where each $\mathrm{op}_{k}$ ($1\leq k\leq m$) is either $\operatorname{OP}_{(i,j),\text{CE}}$ or $\operatorname{OP}_{(i,j),\text{ER}}$ for some $(i,j)$ that can depend on $k$. Because the v-structures from two different sets of the partition have no common vertex by construction and this property is maintained by both operations, we conclude that
$m_{\text{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t}))=\sum_{l=1}^{L}m_{\text{op}}(\mathrm{S}_{l}),$
which lower bounds the number of steps taken to align the v-structures between $\mathcal{G}_{s}$ and $\mathcal{G}_{t}$.
Proposition 6 (CPDAG $d_{\mathcal{L}}$ bounds).
For CPDAGs $\mathcal{G}_{s},\mathcal{G}_{t}\in\mathfrak{G}$, we have upper bound
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\leq\min\left\{d_{\mathcal{L},\downarrow\uparrow}(\mathcal{G}_{s},\mathcal{G}_{t}),d_{\mathcal{L},\uparrow\downarrow}(\mathcal{G}_{s},\mathcal{G}_{t})\right\},$
where $d_{\mathcal{L},\downarrow\uparrow},d_{\mathcal{L},\uparrow\downarrow}$ can be computed using Proposition˜3. Meanwhile, we have lower bound
$d_{\mathcal{L}}(\mathcal{G}_{s},\mathcal{G}_{t})\geq|\operatorname{sk}(\mathcal{G}_{s})\triangle\operatorname{sk}(\mathcal{G}_{t})|+2\,m_{\mathrm{op}}(\mathrm{S}(\mathcal{G}_{s},\mathcal{G}_{t})).$
The upper bound follows from the definition of $d_{\mathcal{L}}$ and the fact that both the down-up and up-down distances are well-defined for CPDAGs, as guaranteed by the existence of the least and the greatest element in its poset. We prove the lower bound in Section˜F.1.